# Mechanistically Interpreting a Transformer-based 2-SAT Solver: An Axiomatic Approach

### v1

@@ Abstract
Mechanistic interpretability aims to reverse engineer the computation performed by a neural network in terms of its internal components. Although there is a growing body of research on mechanistic interpretation of neural networks, the notion of a mechanistic interpretation itself is often ad-hoc. Inspired by the notion of abstract interpretation from the program analysis literature that aims to develop approximate semantics for programs, we give a set of axioms that formally characterize a mechanistic interpretation as a description that approximately captures the semantics of the neural network under analysis in a compositional manner. We use these axioms to guide the mechanistic interpretability analysis of a Transformer-based model trained to solve the well-known 2-SAT problem. We are able to reverse engineer the algorithm learned by the model—the model first parses the input formulas and then evaluates their satisfiability via enumeration of different possible valuations of the Boolean input variables. We also present evidence to support that the mechanistic interpretation of the analyzed model indeed satisfies the stated axioms [^1] Our implementation can be found at https://github.com/nilspalumbo/sat-mi..
@@ 1 Introduction
Neural networks are notoriously opaque. Mechanistic interpretability seeks to reverse-engineer the computation described by a neural network in a human-interpretable form. Typically, the resulting mechanistic interpretation takes the form of a circuit (i.e., program) operating over features (i.e., symbols). The features refer to human-interpretable properties of the input (such as edges, textures, or shapes for vision models and part-of-speech tags, named entities, or semantic relationships for language models) that the model represents internally in its representations spaces whereas the circuit refers to a chain of human-interpretable operations, encoded by the model in its layers, such that each operation processes and transforms the features [39]. While mechanistic interpretability has been used to analyze aspects of vision as well as language models, it has found most success in analyzing small models trained to solve specific algorithmic tasks such as modular addition [40]. Although such small models are not necessarily of practical use, analyzing them has been fruitful for developing the techniques needed to mechanistically interpret.
@@
Despite the broad interest in mechanistic interpretability, we observe that the notion of a mechanistic interpretation has so far only been defined in an ad-hoc manner. In other words, if an analyst claims to have a mechanistic interpretation, there is no agreed-upon standard to judge the analyst’s claims. In this work, we propose a collection of axioms that characterize a mechanistic interpretation. [^2] In the same sense that group axioms characterize the notion of a group in abstract algebra. Our view on mechanistic interpretability is inspired by the notion of abstract interpretation from the program analysis literature [15, 16] that seeks to approximate the semantics of programs in order to make their analysis computationally feasible. Similarly, our axioms are meant to capture the property that the mechanistic interpretation should well approximate the semantics of the neural network under analysis. Further, our axioms also capture the intuition that such an approximation should respect the compositional structure of the neural network—not only should the input-output behaviour of the mechanistic interpretation be as close as possible to the original model, but every step in the reverse-engineered algorithm should be as close as possible to the computation performed by the internal components of the model (i.e., neurons and layers). Our axioms clarify the responsibilities of an analyst; in addition to providing a mechanistic interpretation, an analyst also needs to present evidence to support the fact that the provided interpretation satisfies these axioms.
@@
We demonstrate our approach to mechanistic interpretability on a small neural network with a Transformer-based architecture [54] trained to solve a Boolean satisfiability problem. We choose this as a challenging yet well-understood problem that serves to highlight new techniques and our axiomatic evaluation criteria for mechanistic interpretability. Further, there are many other problems in computer science that reduce to it. To keep the analysis tractable, we focus on the 2-SAT problem with a fixed number of clauses and variables; the problem can be solved in polynomial time (while the more general 3-SAT problem is NP-complete).
@@
Through our analysis, we are able to reverse engineer the algorithm learned by the model—the model parses the formula into a clause-level representation in its initial layers and then uses the later layers to evaluate the formula satisfiability via enumeration of different possible valuations of the Boolean variables. Since we only consider 2-SAT formulas with ten clauses and five variables, such an enumerative approach is computationally feasible and our model has the capacity to encode it. We use novel variants of attention pattern analysis to interpret the first Transformer layer of the model. For the second (i.e., final) Transformer layer, attention pattern analysis does not suffice and we use automated learning of functions (decision trees) to derive an interpretation. We also present evidence that the mechanistic interpretation extracted for this model indeed satisfies our axioms.
@@ 2 Related Work
Work on interpreting neural networks can be broadly divided between techniques that aim to find the input features with the largest effect on model behavior (Input Interpretability) such as gradient-based [50, 52, 32, 51] and activation-based [42, 44, 23, 49, 57] attributions, and those that aim to interpret the internal reasoning performed by a model (Internal Interpretability). We focus on the latter.
@@
Mechanistic Intepretability. In this work, we use mechanistic interpretability to refer to analyses that seek to understand the model behavior completely, i.e., they seek to reverse engineer, in a human-understandable, and hence simplified form, the full algorithm that the neural network implements. Such analyses have typically focused on analyzing toy models trained for algorithmic tasks such as modular addition [40, 67], finding greatest common divisors [10], n-digit integer addition [47], and finite group operations [12]. The standards for judging the validity of the mechanistic interpretations produced by such analyses are ad-hoc and do not take the compositional structure of the neural network into account. We axiomatize the standards for making such judgment in this paper.
Mechanistic interpretability has also been used to refer to analyses that only seek to understand specific aspects of the model behavior. Such analyses, often referred to as circuit analyses, isolate circuits, i.e., subgraphs of the neural network computational graph, that are responsible for the behavior of interest [9, 43, 58, 33, 63, 35, 14]. The circuits are validated by measuring the causal effect of ablating the circuit using techniques such as activation patching [56, 24, 28] on a metric that measures the relevant model behavior. Such analyses are not a subject of our study.
@@
Related to circuit analysis is emerging work on causal abstraction analysis [24, 25, 62, 63] that extracts causal models to explain certain aspects of neural network behavior and uses techniques similar to activation patching for validating the causal abstractions.
Probing. Probing involves training a separate supervised classifier to predict human-understandable features of the data from the model’s internal activations [2, 68] and has found wide applications for understanding the representations learnt by language models [53, 34, 41, 8, 6, 18, 26, 59]. Even though a probe might be successful in predicting features from internal activations, it need not imply a causal relationship between these features and the model output [5]. It has also been questioned whether the model activations genuinely represent the feature of interest or the probe itself learns these features by picking up on spurious correlations [29]. However, recent works have demonstrated that for Transformer-based language models probes can be used to steer the model’s output behavior [68, 34].
Related to probing is the work on concept-based reasoning [31, 17, 65, 64] that extracts representations of high-level concepts as geometric shapes in the neural network representation space (i.e., the activation spaces of inner layers of neural networks) and quantifies the impact of these high-level concepts on model outputs via attribution-based techniques.
Attention Pattern Analysis. There is a large body of work that attempts to understand the behavior of Transformer-based models by analyzing the attention patterns computed by the self-attention layer [3, 66, 21, 55, 27, 46, 38, 1, 11, 37]. However, there is an active ongoing debate about the validity of attention patterns as a tool for interpreting model behavior [30, 45, 7, 4, 61].
@@
The surveys by Millière and Buckner [39] and Räuker et al. [48] are excellent resources for a broader and more detailed description of techniques for interpreting the internals of neural networks.
3 Axioms of Mechanistic Interpretation
@@
At the very least, the input-output behavior of $h$ should be similar to $t$. Ideally, the input-output behaviors of the two programs should coincide on every input but this requirement is much too strong in practice. Assuming that the inputs are drawn from a distribution $\mathcal{D}$, we formalize the weaker requirement that the outputs of the two programs should be equal with a high probability as an axiom. We additionally require that the behaviors of $h$ and $t$ are equivalent at the level of individual components (obtained via suitably decomposing these programs) and that replacing a component of $t$ with the corresponding component of $h$ has a negligible effect on the neural network’s output.
@@
Note that the definition is parametric in decomposition $d_{t}$ of $t$ as well as the input distribution $\mathcal{D}$. In practice, we will often consider layer-wise decompositions of $t$.
Abstraction and Concretization. The first four axioms use the notion of $\alpha$ and $\gamma$ functions which we describe here. The intuition is that a neural network $t$ and a corresponding mechanistic interpretation $h$ operate on different types of data representations. While $t$ operates over real-valued vectors, to aid interpretability, $h$ is intended to operate over human-interpretable features or symbols. Accordingly, $\alpha$ (or abstraction) functions map the real-valued activations (which we also refer to concrete representations) computed by $t$ to the corresponding features or symbols (which we refer to abstract representations) that $h$ operates over. Given a decomposition $d_{t}$ of $t$, we use $\alpha_{i}$ to refer to the $\alpha$ function mapping concrete representations computed by $d_{t}[:i+1]$, i.e., the prefix of length $i$. The $\alpha$ functions are similar to probes [2] for extracting features from a model’s internal activations. The $\gamma$ (or concretization) functions map in the opposite direction—they map abstract features or symbols operated on by $h$ to corresponding real-valued representations of those features in $t$’s representation space. The $\alpha$ and $\gamma$ functions in our axioms are directly inspired by the abstraction and concretization operations from the abstract interpretation literature used to map values between the original semantics and the abstract semantics of a program [15, 16].
Axioms. Axiom 1 bounds the probability that the abstract and the concrete representations computed by the same-length prefix of $h$ and $t$, respectively, do not coincide (after applying $\alpha$ functions to map between the representations). Put differently, the mechanistic interpretation prefixes do not introduce too much error.
Axiom 1 (Prefix Equivalence).
$\forall i\in[len(d)].\underset{x\sim\mathcal{D}}{Pr}[\alpha_{i}\circ d_{t}[i]\circ d_{t}[:i](x)=d_{h}[i]\circ d_{h}[:i]\circ\alpha_{0}(x)]\geq 1-\epsilon$
In contrast, Axiom 2 requires that none of the individual components of the mechanistic interpretation $h$ introduce too much error. Axiom 2 does not imply Axiom 1 since the errors introduced by each component of the mechanistic interpretation can compound in the worst case [20].
Axiom 2 (Component Equivalence).
$\forall i\in[len(d)].\underset{x\sim\mathcal{D}}{Pr}[\alpha_{i}\circ d_{t}[i]\circ d_{t}[:i](x)=d_{h}[i]\circ\alpha_{i-1}\circ d_{t}[:i](x)]\geq 1-\epsilon$
@@
Axiom 3 (Prefix Replaceability).
$\forall i\in[len(d)].\underset{x\sim\mathcal{D}}{Pr}[t(x)=d_{t}[i+1:]\circ\gamma_{i}\circ d_{h}[i]\circ d_{h}[:i]\circ\alpha_{0}(x)]\geq 1-\epsilon$
Axiom 4 (Component Replaceability).
$\forall i\in[len(d)].\underset{x\sim\mathcal{D}}{Pr}[t(x)=d_{t}[i+1:]\circ\gamma_{i}\circ d_{h}[i]\circ\alpha_{i-1}\circ d_{t}[:i](x)]\geq 1-\epsilon$
@@
We propose two additional axioms, namely, Axioms 5 and 6, that are more informal in nature and we do not consider them for the analysis presented in this paper. These axioms are presented as a goal for future work and described in Appendix A.
@@
Checking the Axioms. The proposed axioms can be checked statistically. Given a test dataset (which need not be labeled), we can estimate the error rate $\epsilon$ for each axiom by simply computing the proportion of inputs on which the conditions specified by the axioms are satisfied. Confidence intervals for our estimates of $\epsilon$ can be computed using any method for the parameter $p$ of a Binomial distribution (Clopper-Pearson method [13] in our experiments). Validating the axioms is thus cheap and feasible in practice. These axioms represent a minimal set that we believe are necessary albeit may not be sufficient in all cases for characterizing a mechanistic interpretation. In future work, we plan to address this by considering analysis of models trained for other problems.
@@
Relationship to Nanda et al. [40]’s Analysis. The seminal work of Nanda et al. [40] on mechanistically interpreting the complete behavior of a toy model serves as an interesting case study to demonstrate the relationship between the existing standards for judging the validity of a mechanistic interpretation and our proposed axioms. Nanda et al. [40] present four different pieces of evidence in support of the validity of their interpretation. The first three pieces of evidence (presented in Sections 4.1, 4.2, and 4.3 in their paper) resemble our Axiom 2 that compares each individual component of the mechanistic interpretation with the corresponding component of the model. The evidence they present in Section 4.4 of their paper is similar to our Axiom 4 that checks the effect on the output of the model when individual model components are replaced by the corresponding components from the mechanistic interpretation. Notably, they do not present the evidence required by our Axioms 1 and 3. This amounts to not considering the compositional structure of the model and, therefore, the possible compounding of error introduced by each component of the mechanistic interpretation.
4 Data and Model Details

### v2

# Validating Mechanistic Interpretations: An Axiomatic Approach

@@ Abstract
Mechanistic interpretability aims to reverse engineer the computation performed by a neural network in terms of its internal components. Although there is a growing body of research on mechanistic interpretation of neural networks, the notion of a mechanistic interpretation itself is often ad-hoc. Inspired by the notion of abstract interpretation from the program analysis literature that aims to develop approximate semantics for programs, we give a set of axioms that formally characterize a mechanistic interpretation as a description that approximately captures the semantics of the neural network under analysis in a compositional manner. We demonstrate the applicability of these axioms for validating mechanistic interpretations on an existing, well-known interpretability study as well as on a new case study involving a Transformer-based model trained to solve the well-known 2-SAT problem.
mechanistic interpretability, 2-SAT solver, axiomatic approach
@@ 1 Introduction
Neural networks are notoriously opaque. Mechanistic interpretability seeks to reverse-engineer the computation described by a neural network in a human-interpretable form. Typically, the resulting mechanistic interpretation takes the form of a circuit (i.e., program) operating over features (i.e., symbols). The features refer to human-interpretable properties of the input (such as edges, textures, or shapes for vision models and part-of-speech tags, named entities, or semantic relationships for language models) that the model represents internally in its representations spaces whereas the circuit refers to a chain of human-interpretable operations, encoded by the model in its layers, such that each operation processes and transforms the features (Millière & Buckner, 2024). While mechanistic interpretability has been used to analyze aspects of vision as well as language models, it has found most success in analyzing small models trained to solve specific algorithmic tasks such as modular addition (Nanda et al., 2023a). Although such small models are not necessarily of practical use, analyzing them has been fruitful for developing the techniques needed to mechanistically interpret.
@@
Despite the broad interest in mechanistic interpretability, we observe that the notion of a mechanistic interpretation has so far been primarily defined in an ad-hoc manner. In this work, we propose a collection of axioms that characterize a mechanistic interpretation. [^1] In the same sense that group axioms characterize the notion of an abstract group in abstract algebra. Our view on mechanistic interpretability is inspired by the notion of abstract interpretation from the program analysis literature (Cousot & Cousot, 1977, 1992) that seeks to approximate the semantics of programs in order to make their analysis computationally feasible. Similarly, our axioms are meant to capture the property that the mechanistic interpretation should capture the semantics of the model. Further, our axioms also capture the intuition that such an approximation should respect the compositional structure of the neural network—not only should the input-output behavior of the mechanistic interpretation be as close as possible to the original model, but every step in the reverse-engineered algorithm should be as close as possible to the computation performed by the internal components of the model (i.e., neurons and layers). Our axioms clarify the responsibilities of an analyst; in addition to providing a mechanistic interpretation, an analyst also needs to present evidence to support the fact that the provided interpretation is valid, i.e., satisfies these axioms. Note that our axioms present a general framework for validating mechanistic interpretations. Closest to our work are recent papers by Geiger et al. (2021) and Chan et al. (2022) that seek to formalize the notion of mechanistic interpretability from a causal perspective. We believe that these formulations are complementary to our axioms.
@@
We demonstrate the applicability of our axioms for validating mechanistic interpretations via two case studies [^2] Our implementation is available at https://github.com/nilspalumbo/axiomatic-validation.—the existing, well-known analysis by Nanda et al. (2023a) of a model with a Transformer-based architecture (Vaswani et al., 2017) trained to solve modular addition, and a new analysis of a small Transformer-based model trained by us to solve a Boolean satisfiability problem. For the new case study, we choose Boolean satisfiability as a challenging yet well-understood problem that serves to highlight new techniques and our axiomatic evaluation criteria for mechanistic interpretability. Further, there are many other problems in computer science that reduce to it, for example, graph reachability is reducible to 2-SAT. To keep the analysis tractable, we focus on the 2-SAT problem with a fixed number of clauses and variables; the problem can be solved in polynomial time (while the more general 3-SAT problem is NP-complete).
@@
To apply our axioms, we first need to extract a mechanistic interpretation of the model under analysis. Although such an interpretation is already known for the modular arithmetic model, we present a new analysis for the 2-SAT model that is guided by our axioms. Through our analysis, we are able to reverse engineer the algorithm learned by the 2-SAT model—the model parses the formula into a clause-level representation in its initial layers and then uses the later layers to evaluate the formula satisfiability via enumeration of different possible valuations of the Boolean variables. Since we only consider 2-SAT formulas with ten clauses and five variables, such an enumerative approach is computationally feasible and our model has the capacity to encode it. We use novel variants of attention pattern analysis to interpret the first Transformer block of the model. For the second (i.e., final) Transformer block, attention pattern analysis does not suffice and we use automated learning of functions (decision trees) to derive an interpretation. Finally, for both the case studies, we present evidence that the mechanistic interpretations extracted for these models indeed satisfy our axioms.
@@ 2 Related Work
Work on interpreting neural networks can be divided between techniques that find input features with the largest effect on model behavior (Input Interpretability) such as gradient-based (Simonyan et al., 2013; Sundararajan et al., 2017; Leino et al., 2018; Smilkov et al., 2017) and activation-based (Olah et al., 2017; Petsiuk et al., 2018; Fong & Vedaldi, 2017; Selvaraju et al., 2019; Wang et al., 2020) attributions, and those that interpret the internal reasoning performed by a model (Internal Interpretability). We focus on the latter.
@@
Mechanistic interpretability typically refers to analyses that seek to understand the model behavior completely, i.e., they seek to reverse engineer, in a human-understandable, and hence simplified form, the full algorithm that the neural network learns. Such analyses have tended to focus on analyzing toy models trained for algorithmic tasks such as modular addition (Nanda et al., 2023a; Zhong et al., 2023), finding greatest common divisors (Charton, 2023), $n$-digit integer addition (Quirke & Barez, 2024), and finite group operations (Chughtai et al., 2023). Mechanistic interpretability has also been used to refer to analyses that only seek to understand specific aspects of the model behavior. Such analyses, often referred to as circuit analyses, isolate circuits, i.e., subgraphs of the neural network computational graph, that are responsible for the behavior of interest (Cammarata et al., 2020; Olsson et al., 2022; Wang et al., 2023; Lepori et al., 2024; Wu et al., 2023b; Lieberum et al., 2023; Conmy et al., 2023). The circuits are validated by measuring the causal effect of ablating the circuit using techniques such as activation patching (Vig et al., 2020; Geiger et al., 2021; Heimersheim & Nanda, 2024) on a metric that measures the relevant model behavior. In either case, the standards for judging the validity of the mechanistic interpretations produced by such analyses are ad-hoc and do not take the compositional structure of the neural network into account. We axiomatize the standards for making such judgment in this paper.
@@
Related to circuit analysis and the evaluation of mechanistic interpretations is emerging work on causal abstraction analysis (Geiger et al., 2021, 2024; Wu et al., 2023a, b; Chan et al., 2022) that extracts causal models to explain certain aspects of neural network behavior and uses techniques similar to activation patching for validating the causal abstractions.
@@
The surveys by Millière & Buckner (2024), Mueller et al. (2024), and Räuker et al. (2023) are excellent resources for a broader and more detailed description of techniques for interpreting the internals of neural networks and the architectural components which form the interpreted units. Appendix A has a further discussion on internal interpretability techniques such as probing and attention pattern analysis.
3 Mechanistic Interpretation Axioms
@@
At the very least, the input-output behavior of $h$ should be similar to $t$. Ideally, the input-output behaviors of the two programs should coincide on every input but this requirement is much too strong in practice. Assuming that the inputs are drawn from a distribution $\mathcal{D}$, we formalize the weaker requirement that the outputs of the two programs should be equal with a high probability as an axiom. For these equality conditions to be practically applicable, $h$ must operate over discrete values. Also, the output type of $t$ must be discrete, in particular, we assume that an $\operatorname{argmax}$ or $\operatorname{top-k}$ operator has been applied to the logits. We additionally require that the behaviors of $h$ and $t$ are equivalent at the level of individual components (obtained via suitably decomposing these programs) and that replacing a component of $t$ with the corresponding component of $h$ has a negligible effect on the neural network’s output.
@@
Note that the definition is parametric in decomposition $d_{t}$ of $t$ as well as the input distribution $\mathcal{D}$. In practice, we will often consider layer or block-wise decompositions of $t$.
Abstraction and Concretization. The first four axioms use the notion of $\alpha$ and $\gamma$ functions which we describe here. The intuition is that a neural network $t$ and a corresponding mechanistic interpretation $h$ operate on different types of data representations. While $t$ operates over real-valued vectors, to aid interpretability, $h$ is intended to operate over human-interpretable features or symbols. Accordingly, $\alpha$ (or abstraction) functions map the real-valued activations (which we also refer to concrete representations) computed by $t$ to the corresponding features or symbols (which we refer to abstract representations) that $h$ operates over. Given a decomposition $d_{t}$ of $t$, we use $\alpha_{i}$ to refer to the $\alpha$ function mapping concrete representations computed by $d_{t}[:i+1]$, i.e., the prefix of length $i$. The $\alpha$ functions are similar to probes (Alain & Bengio, 2017) for extracting features from a model’s internal activations. The $\gamma$ (or concretization) functions map in the opposite direction—they map abstract features or symbols operated on by $h$ to corresponding real-valued representations of those features in $t$’s representation space. The $\alpha$ and $\gamma$ functions in our axioms are directly inspired by the abstraction and concretization operations from the abstract interpretation literature used to map values between the original semantics and the abstract semantics of a program (Cousot & Cousot, 1977, 1992). We note that the $\alpha$ and $\gamma$ functions need to be individually instantiated every mechanistic interpretation.
Axioms. We define the following axioms for a particular choice of $\epsilon$; in our analysis we say that an interpretation satisfies an axiom with $\epsilon$ when the statement of that axiom holds.
Axiom 1 bounds the probability that the abstract and the concrete representations computed by the same-length prefix of $h$ and $t$, respectively, do not coincide (after applying $\alpha$ functions to map between the representations). Put differently, the mechanistic interpretation prefixes do not introduce too much error.
Axiom 1 ($\epsilon$-Prefix Equivalence).
$\forall i\in[len(d)].$
$\underset{x\sim\mathcal{D}}{Pr}[\alpha_{i}\circ d_{t}[:i+1](x)=d_{h}[:i+1]\circ\alpha_{0}(x)]\geq 1-\epsilon$
In contrast, Axiom 2 requires that none of the individual components of the mechanistic interpretation $h$ introduce too much error. Axiom 2 does not imply Axiom 1 since the errors introduced by each component of the mechanistic interpretation can compound in the worst case (Dziri et al., 2023).
Axiom 2 ($\epsilon$-Component Equivalence).
$\forall i\in[len(d)].$
$\underset{x\sim\mathcal{D}}{Pr}[\alpha_{i}\circ d_{t}[:i+1](x)=d_{h}[i]\circ\alpha_{i-1}\circ d_{t}[:i](x)]\geq 1-\epsilon$
@@
Axiom 3 ($\epsilon$-Prefix Replaceability).
$\forall i\in[len(d)].$
$\underset{x\sim\mathcal{D}}{Pr}[t(x)=d_{t}[i+1:]\circ\gamma_{i}\circ d_{h}[:i+1]\circ\alpha_{0}(x)]\geq 1-\epsilon$
Axiom 4 ($\epsilon$-Component Replaceability).
$\forall i\in[len(d)].$
$\underset{x\sim\mathcal{D}}{Pr}[t(x)=d_{t}[i+1:]\circ\gamma_{i}\circ d_{h}[i]\circ\alpha_{i-1}\circ d_{t}[:i](x)]\geq 1-\epsilon$
@@
We propose two additional axioms, namely, Axioms 5 and 6, that are more informal in nature and we do not consider them for the analysis presented in this paper. These axioms are presented as a goal for future work and described in Appendix B.
@@
Checking the Axioms. The proposed axioms can be checked statistically. Given a test dataset (which need not be labeled), we can estimate the error rate $\epsilon$ for each axiom by computing the proportion of test inputs which violate the equality condition specified by the axiom. The number of violations is distributed binomially—the parameter $p$ of this binomial distribution is equal to $\epsilon$. Hence, any method for deriving confidence intervals for the parameter of a binomial distribution can be used to derive confidence intervals for our estimate of $\epsilon$; we use the Clopper-Pearson method (Clopper & Pearson, 1934) in our experiments. Validating the axioms is thus cheap and feasible in practice. These axioms represent a minimal set that we believe are necessary albeit may not be sufficient in all cases for characterizing a mechanistic interpretation. In future work, we plan to address this by considering analysis of models trained for other problems.
Relationship to Existing Approaches. The evaluation of Nanda et al. (2023a)’s seminal paper on mechanistically interpreting the complete behavior exemplifies typical approaches to evaluating a mechanistic interpretation (Zhong et al., 2023; Quirke & Barez, 2024; Chughtai et al., 2023). The first three pieces of evidence (presented in Sections 4.1, 4.2, and 4.3 in their paper) resemble our Axiom 2 that compares each individual component of the mechanistic interpretation with the corresponding component of the model. However, as there does not exist a direct analogue of abstraction, this analysis is primarily observational.
@@
The evidence they present in Section 4.4 of their paper is similar to our Axiom 4 that checks the effect on the output of the model when individual model components are replaced by the corresponding components from the mechanistic interpretation. Notably, they do not present the evidence required by our Axioms 1 and 3. This amounts to not considering the compositional structure of the model and, therefore, the possible compounding of error introduced by each component of the mechanistic interpretation; we discuss this further at the end of this section.
Causal abstraction (Geiger et al., 2021, 2022, 2024, 2025) and causal scrubbing (Chan et al., 2022) are key existing attempts to formalize the notion of a mechanistic interpretation. Both of these formulate the abstract and concrete models as directed acyclic graphs (DAG) with a map that relates the two DAGs. While it may seem that, compared to our approach, the DAG formulation is more generic since it admits parallel composition of components in addition to sequential composition, this is not fundamental, in particular, Appendix B.1 discusses how to represent arbitrary computational graphs in our framework.
The map used by causal abstraction to relate the two DAGs corresponds to our abstraction function $\alpha$; however, it does not have any analog of a concretization function $\gamma$. Geiger et al. (2025) note that the inverse of abstraction may be used as a concretization operator; however, this may not be feasible in practice, as this choice necessitates the use of set semantics for evaluation of the concrete model when the abstraction operator fails to be invertible. Under the causal abstraction framework, a valid interpretation is defined by comparing the effects of interventions on the concrete model with corresponding interventions on the abstract model; in contrast, our notion of a valid interpretation may be viewed as one which is invariant up to a class of interventions, characterized by our axioms, which interleave the concrete and abstract models. While causal abstraction is a very general framework for the validation of mechanistic interpretations, in practice, it uses the specific metric of interchange intervention accuracy (Geiger et al., 2022) to validate interpretations and this metric fails to directly evaluate the equivalence of internal representations.
On the other hand, the map defined by causal scrubbing can be seen as an form of concretization, but it does not define an abstraction function. While both these approaches compare the outputs of the concrete and abstract models similar to our Axioms 3 and 4, they tend to omit direct evaluation of the equivalence of internal representations, as in our Axioms 1 and 2. This is potentially problematic since there is no direct validation that the intermediate steps of the mechanistic interpretation correspond to the intermediate steps of the neural network (Scheurer et al., 2023). Further, as causal abstraction and scrubbing lack concretization and abstraction operators respectively, even Axioms 3 and 4 cannot be expressed directly in these frameworks.
Compositionality. While it may appear that our axioms for compositional evaluation (Axioms 1 and 3) follow from their componentwise counterparts (Axioms 2 and 4 respectively), this is not the case; for example, while bounds of the form of Axiom 1 can be derived from Axiom 2, these hold only with far weaker $\epsilon$; when the number of components is sufficiently high, the derived bounds entirely cease to be meaningful. More details are given in Appendix E; Appendix F includes empirical evidence demonstrating the pitfalls of omitting the compositional axioms.
4 Case Study: 2-SAT
4.1 Data and Model Details